A store buys jeans for 24 dollars each and marks them up 30 percent what's the selling price of a pair of jeans
Answer by ella13(7)

$24*0.3=7.2
24+7.2=31.2$
% has to be change to decimal 30/100=0.3 then add to the jeans price to get the total price.